OpenStudy (anonymous):

Which of the following is a solution to 4sin^2x – 1 = 0? A.30 B.60 C.90 D.120

OpenStudy (anonymous):

do you know what the value of sin30, sin60, sin90, and sin120 ?

OpenStudy (anonymous):

sin30=\[\sqrt{3}\div2\], sin60=1/2, sin90=0, sin120=-1/2

OpenStudy (anonymous):

huh ? i knowed that is for cos values :)

OpenStudy (anonymous):

sorrry i confused myself its sin30=1/2, sin60=\[\sqrt{3}\div2\],sin90=1, and sin120= the same for sin60

OpenStudy (anonymous):

yes, that's look better :) so, just check one by one in your answer choice. example : (this is for wrong answer, let it is C) 4sin^2x – 1 = 0 4sin^2(90) – 1 = 0 4(1)^2 - 1 = 0 4 - 1 = 0 3 = 0 LHS not equal with RHS

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)